Find the number of 4-letter code words that can be formed from the letters of the word "advance".
1.using neither a.
2.both a's

Answer by Edwin McCravy(20055)

Find the number of 4-letter code words that can be formed from the letters of the word "advance".
1.using neither a.
Then we place four letters only from the set {d,v,n,c,e}

Answer:  P(5,4) = 5×4×3×2 = 120

2.both a's
We choose the two positions {1st,2nd,3rd,4th} for the two a's 
That's C(4,2) ways.
We place the other two letters P(5,2) ways.

Answer: C(4,2)P(5,2) = 6×20 = 120

Edwin